OpenStudy (xenesis98):

MEDALS!!Which expression is equivalent to tan(-x)cos^2x? a.-sin xcos x b.sin xcos x c.tan x + sin^2x d.tan x − sin^2x e.-tan x − sin^2x

OpenStudy (anonymous):

a.

OpenStudy (anonymous):

\[\tan \left( -x \right)=-\tan x=-\frac{ \sin x }{ \cos x }\]

OpenStudy (anonymous):

\[\tan (-x)=-\tan (x) \]\[-\frac{\sin (x) \cos (x) \cos (x)}{\cos (x)}=-\sin (x) \cos (x) \]